# Algebraic Independence of Special Points on Shimura Varieties

### v1

@@ Abstract.
Given a correspondence $V$ between a connected Shimura variety $S$, a commutative connected algebraic group $T$, and $n\in\mathbb{N}$, we prove that the $V$-images of any $n$ special points on $S$ outside a proper Zariski closed subset are algebraically independent. Our result unifies previous unlikely intersection results on multiplicative independence and linear independence. We prove multiplicative independence of differences of singular moduli, generalizing previous results by Pila-Tsimerman, and Aslanlyan-Eterović-Fowler. We also give an application to abelian varieties by proving that the special points of $S$ whose $V$-images lie in a finite-rank subgroup of $T$ are contained in a finite union of proper special subvarieties of $S$, only dependent on the rank of the subgroup. In this way, our result is a generalization of the works of Pila-Tsimerman and Buium-Poonen.
@@ 1.1. Statement of the results
Let $S$ be a connected Shimura variety, $T$ be a commutative connected algebraic group. Let $V\subset S\times T$ be a proper irreducible closed subvariety such that the projection on each factor is dominant, and $V$ is finite over $S$. Given a point $(s,t)\in V\subset S\times T$, we call $t$ a $V$-image of $s$.
The special subvarieties in $T$ are torsion cosets of connected algebraic subgroups and the special subvarieties in a Shimura variety are described in Section 2.1. We prove the independence of the $V$-images of special points of $S$. Previous work in the literature proved the linear independence of special points when $V$ is a correspondence from a (product of) modular curve or a Shimura curve to an elliptic curve, i.e. the linear dependences among Heegner points, which plays an important role in the study of the Birch and Swinnerton-Dyer conjecture. See [BP09], [Bal18], [PT22] for results in this direction. There have also been results proving multiplicative independence of special points when $V$ is the graph of the $j$-function from $Y(1)\to\mathbb{G}_{m}$ (see [PT17], [AEF23]). The aim of this article is to provide a unified framework to generalize the previous works to the study of algebraic independence of special points in general commutative groups.
@@
With the notation as above, let $\pi_{S}$ and $\pi_{T}$ be the projections of $S\times T$ onto the first and second factors respectively.
  • A distinguished component is an irreducible component of $W\subset V\cap(S^{\prime}\times T^{\prime})$, where $S^{\prime}\times T^{\prime}$ is a special subvariety of $S\times T$, such that $\pi_{S}(W)=S^{\prime}$.
  • Let $W$ be a distinguished component and let $B\supset\pi_{T}(W)$ be the smallest special subvariety of $T$ containing $\pi_{T}(W)$. We say that $W$ is exemplary if there are no larger distinguished components $W^{\prime}\supset W$ such that $\pi_{T}(W^{\prime})\subset B$.
  • We say a distinguished component $W$ in $T$ is dependent if $T^{\prime}\subset T$ is strict. We note that the unique non-dependent exemplary component is $V$ itself.
Pila and Tsimerman [PT22, Theorem 1.1] were able to deal with the case where $S=Y(1)^{n}$ is the $n$-copy of a modular (or Shimura) curve and $T=E^{n}$ is the self-$n$-product of an elliptic curve for $n\geq 1$. We prove the following theorem, which is a generalization of their main theorem to arbitrary (connected) Shimura varieties and commutative groups. The theorem describes all algebraic dependence among $V$-images of $n$ special points for any $n\geq 1$.
@@
Let $S$ be a connected Shimura variety, $T$ be a commutative connected algebraic group. Suppose $V\subset S\times T$ is an irreducible subvariety that maps finitely to $S$. Then, there are only finitely many exemplary components in $V$.
@@
Let $S,T,V$ be as above. Suppose moreover that the projections of $V$ to $S$ and $T$ are both dominant. Fix $n\in\mathbb{N}$. There exists a proper Zariski closed subset $S^{\prime}\subset S^{n}$ such that for any $n$ points $s_{1},\dots,s_{n}\in S$ satisfying $(s_{1},\dots,s_{n})\not\in S^{\prime}$, any $V$-images $t_{1},\dots,t_{n}$ of $s_{1},\dots,s_{n}$ are linearly independent.
@@
For $n\geq 1$, let $\boldsymbol{s}=(s_{1},\cdots,s_{n})$ be a special point in $S^{n}$ and let $\boldsymbol{t}=(t_{1},\cdots,t_{n})$ be a point in $T^{n}$ such that $(\boldsymbol{s},\boldsymbol{t})\in V^{n}$. It follows from the definition that if $(\boldsymbol{s},\boldsymbol{t})\in W$ for some dependent exemplary component then $t_{1},\cdots,t_{n}$ are algebraically dependent in $T$. Conversely, if $t_{1},\cdots,t_{n}$ are algebraically dependent in $T$, then $(\boldsymbol{s},\boldsymbol{t})$ is a distinguished point and hence contained in some dependent exemplary component. ∎
@@
Suppose that $S\subset\mathcal{A}_{g}$ for $g\leq 3$. For $n\geq 1$ there exists a positive integer $D=D(n,S,T,V)$ such that if $s_{1},\dots,s_{n}$ are $D$-independent, then any $V$-image $x_{1},\dots,x_{n}$ of $s_{1},\dots,s_{n}$ are algebraically independent in $T$.
Taking $V$ to be the graph of the function $f\colon Y(1)^{n}\times Y(1)^{2n}\to\mathbb{C}^{n}$ given by
@@
For $n\geq 1$, there exists a positive integer $D=D(n)$ such that if $x_{1},\dots,x_{n},$ $y_{1},\dots,y_{n}\in Y(1)$ are singular moduli that are $D$-independent, then there do not exist $\0a_{1},\dots,a_{n}\in\mathbb{Z}\backslash\{0\}$ such that
@@
We prove also that on low dimensional abelian varieties, the intersection of any finite-rank subgroup of $T$ with the set of $V$-images of special points on $S$ is contained in a finite union of proper special subvarieties. This generalizes Pila and Tsimerman’s result [PT22, Corollary 1.4] to higher dimensional abelian varieties, and Buium and Poonen’s [BP09, Theorem 1.1] in a uniform way, depending only on the rank of the subgroup.
@@
Suppose $T$ is defined over a field $K$ of characteristic 00 and suppose the Shimura datum for $S$ has $G$ simple. Let $\Gamma\subset T(K)$ be a subgroup of rank $r$ and let $\Gamma^{\prime}$ be its division group. Then, there exist $N(r,S,T,V)$ proper special subvarieties of $S$ such that the set of special points of $S$ with a $V$-image in $\Gamma^{\prime}$ is contained in a union of these proper special subvarieties.
A similar argument to [PT22, Proposition 3.2] proves that Theorem 1.2 is a consequence of the Zilber–Pink Conjecture. Also, Theorem 1.6 may be thought as a weak version of the André–Oort–Mordell–Lang conjecture in the case of subgroups of finite rank formulated by Baldi in [Bal18, Conjecture 5.2]. The original formulation of the André–Oort–Mordell–Lang conjecture is due to Pink which can be found in [Pin05].
@@ 1.2. Organization of the paper
The notions of Shimura varieties, special and weakly special subvarieties are explained in Section 2. We also proved a mixed version of the Ax–Schanuel theorem for the uniformization of the product $S\times T$ in the form needed for our main theorem in this section. In Section 3, we collect and prove some arithmetic estimates. The main theorem is proved in Section 4 and its applications and corollaries are proved in Section 5.
@@ 2. Shimura varieties and the mixed Ax–Schanuel
Let $S$ be an algebraic variety and $q\colon X\to S$ be a universal covering of $S$. The Ax–Schanuel theorem gives information about the bi-algebraic varieties, algebraic varieties $V\subset X$ such that $q(V)$ is also algebraic, for the transcendental covering map $q$. The rough statement of the Ax–Schanuel theorem is that varieties of $X\times S$ should have a proper intersection with the graph of $q$ unless the projection of $V$ is contained within a bi-algebraic variety. These bi-algebraic varieties are precisely the weakly special subvarieties. We first give precise definitions of bi-algebraic varieties and results in the case when $S$ is a semi-abelian variety or a Shimura variety, and then prove the result when $S$ is a product of a semi-abelian variety with a Shimura variety.
@@ 2.1. Shimura Varieties
Let $S:=\operatorname{Sh}_{K}(G,X)$ be the (connected) Shimura variety associated with a connected Shimura datum $(G,X)$, where $G$ is a semisimple group of adjoint type and $K$ is a compact open subgroup of $G(\mathbb{A}_{f}).$ More information about Shimura varieties can be found in Milne ([Mil05]).
@@
For any $x\in X$, let $\mathbf{M}:=\operatorname{MT}(x)$ be the Mumford-Tate group of $x$, which is defined as the smallest $\mathbb{Q}$-subgroup $H$ of $G$ such that $x$ factors through $H_{\mathbb{R}}$. Let $X_{\mathbf{M}}$ be the $\mathbf{M}(\mathbb{R})$ conjugacy class of $x$. Then the image of $X_{\mathbf{M}}$ in $\Gamma_{g}\backslash X$ is a special subvariety of $S$. It is not hard to see that every special subvariety of $S$ arises this way. Since the action of $\mathbf{M}(\mathbb{R})$ on $X_{\mathbf{M}}$ factor through the adjoint group $\mathbf{M}^{\operatorname{ad}}(\mathbb{R})$, which is a direct product of its $\mathbb{Q}$-simple factors. Therefore we can write $X_{\mathbf{M}}$ as a product
$X_{\mathbf{M}}=X_{1}\times X_{2}$
corresponds to the action of $M^{\operatorname{ad}}(\mathbb{R})$. The weakly special subvariety is the image of the fiber $\{x_{1}\}\times X_{2}$ or $X_{1}\times\{x_{2}\}$ for any $x_{1}\in X_{1}$ or $x_{2}\in X_{2}$. Therefore, a weakly special subvariety of $S$ is a special subvariety if and only if it contains a special point. The weakly special subvarieties of $S$ are precisely those subvarieties that are totally geodesic in $S$ ([Moo98, Section 4]) and they are also precisely the bi-algebraic subvarieties under the uniformization map $q\colon X\to S$. More detailed information of weakly special subvarieties of Shimura varieties can be found in [UY11].
@@
Then the projection of $U$ to $S$ is contained in a proper weakly special subvariety of $S$.
2.2. Semi-abelian Varieties
Let $T$ be an semi-abelian variety of dimension $g$ defined over $\mathbb{C}$ and let $\exp\colon\mathbb{C}^{g}\to T$ be the exponential map from $\operatorname{Lie}T\cong\mathbb{C}^{g}$ to $T$.
@@ 2.2. Semi-abelian Varieties
For each semi-abelian subvariety $B\subset T$ and closed point $p\in T(\mathbb{C})$, we say that the algebraic subvariety $B+p\subset T$ is a weakly special subvariety of $T$. We say that it is a special subvariety of $T$ if $p\in T_{\mathrm{tors}}$ is a torsion point.
The special points of $T$ are special subvarieties of dimension 00, and hence they are precisely the torsion points of $T$. As before, a weakly special subvariety of $T$ is special precisely when it contains a special point.
The Ax–Schanuel theorem for semi-abelian varieties is a statement about covering map given by the $\operatorname{Lie}$-exponential $\exp$.
@@
Let $V\subset\mathbb{C}^{g}\times T$ be an algebraic subvariety and let $D\subset\mathbb{C}^{g}\times T$ be the graph of $\exp\colon\mathbb{C}^{g}\to T$. Let $U\subset V\cap D$ be an irreducible component such that
$\dim U>\dim V-\dim T.$
Then the projection of $U$ to $T$ is contained in a proper weakly special subvariety of $T$.
@@ 2.3. Mixed Ax–Schanuel
We will prove a version of the Ax–Schanuel theorem for the product uniformization map $q\times\exp\colon X\times\mathbb{C}^{g}\to S\times T$. This was proven by Pila and Tsimerman in [PT22] for the case when $S$ is a product of modular curves and $T$ is a product of elliptic curves. It was shown that the Ax–Schanuel theorem for the product follows from an Ax–Schanuel theorem for mixed Shimura varieties by Gao in [Gao20]. Using the same method, one can prove that the Ax–Schanuel theorem holds for the product $S\times T$ whenever $S$ is a Shimura variety of abelian type. However, to prove it for general Shimura varieties and general semi-abelian varieties, we will need stronger machinery from Blázquez-Sanz, Casale, Freitag, and Nagloo ([BSCFN23]). The translation into the language of (weakly) special subvarieties is given by Chiu ([Chi22]).
@@
Let $S$ be a Shimura variety and $T$ a semi-abelian variety. We say that an algebraic subvariety $V\subset S\times T$ is a weakly special subvariety of $S\times T$ if there exist weakly special subvarieties $S_{H}\subset S$ and $B\subset T$ such that $V=S_{H}\times B$. We say that $V$ is a special subvariety if $S_{H},B$ are both special.
@@
Let $V\subset(X\times\mathbb{C}^{g})\times(S\times T)$ be an algebraic subvariety and let $D\subset(X\times\mathbb{C}^{g})\times(S\times T)$ be the graph of $q\times\exp\colon(X\times\mathbb{C}^{g})\to(S\times T)$. Let $U\subset V\cap D$ be an irreducible component such that
$\dim U>\dim V-\dim S\times T.$
Then the projection of $U$ to $S\times T$ is contained in a proper weakly special subvariety of $S\times T$.
@@
Let the Shimura datum of $S$ be $(G,X)$. We may take $S$ to be a connected Shimura variety and $G$ to be a derived group. Applying [BSCFN23, Thm. 3.6] to $D\subset(X\times\mathbb{C}^{g})\times(S\times T)$ gives that the projection of $U$ in $S\times T$ is contained in a proper subvariety whose Galois group is a strict algebraic subgroup $H\subset G(\mathbb{C})\times\mathbb{G}_{a}(\mathbb{C})^{g}$. Since $G$ is a derived group, it has no abelian quotients and hence Goursat’s lemma says that the projection of $H$ to $G(\mathbb{C})$ or $\mathbb{G}_{a}(\mathbb{C})^{g}$ is not onto. By [Chi22, Thm. 3.2], that means the projection of $U$ to $S$ or $T$ is contained in a proper weakly special subvariety. ∎

### v2

@@ Abstract.
Given a correspondence $V$ between a connected Shimura variety $S$ and a commutative connected algebraic group $G$, and $n\in\mathbb{N}$, we prove that the $V$-images of any $n$ special points on $S$ outside a proper Zariski closed subset are algebraically independent. Our result unifies previous unlikely intersection results on multiplicative independence and linear independence. We prove multiplicative independence of differences of singular moduli, generalizing previous results by Pila–Tsimerman, and Aslanyan–Eterović–Fowler. We also give an application to abelian varieties by proving that the special points of $S$ whose $V$-images lie in a finite-rank subgroup of $G$ are contained in a finite union of proper special subvarieties of $S$, only dependent on the rank of the subgroup. In this way, our result is a generalization of the works of Pila–Tsimerman and Buium–Poonen.
@@ 1.1. Statement of the results
Let $S$ be a connected Shimura variety, $G$ be a commutative connected algebraic group. Let $V\subset S\times G$ be a proper irreducible closed subvariety such that the projection on each factor is dominant, and $V$ is finite over $S$. Given a point $(s,g)\in V\subset S\times G$, we call $g$ a $V$-image of $s$. Moreover, we say that a series of points $g_{1},\dots,g_{n}\in G$ are algebraically independent if they do not all lie in a proper algebraic subgroup.
The special subvarieties in $G$ are torsion cosets of connected algebraic subgroups and the special subvarieties in a Shimura variety are described in Section 2.A. We prove the independence of the $V$-images of special points of $S$. Previous work in the literature proved the linear independence of special points when $V$ is a correspondence from a modular curve or a Shimura curve to an elliptic curve, i.e. the linear dependences among Heegner points, which play an important role in the study of the Birch and Swinnerton-Dyer Conjecture. See [BP09], [Bal18], [PT22] for results in this direction. There have also been results proving multiplicative independence of special points when $V$ is the graph of the $j$-function from $Y(1)\to\mathbb{G}_{m}$ (see [PT17], [AEF23]). The aim of this article is to provide a unified framework to generalize the previous works to the study of algebraic independence of special points in general commutative groups.
@@
With the notation as above, let $\pi_{S}$ and $\pi_{G}$ be the projections of $S\times G$ onto the first and second factors respectively.
  • A distinguished component is an irreducible component of $W\subset V\cap(S^{\prime}\times G^{\prime})$, where $S^{\prime}\times G^{\prime}$ is a special subvariety of $S\times G$, such that $\pi_{S}(W)=S^{\prime}$.
  • Let $W$ be a distinguished component and let $B\supset\pi_{G}(W)$ be the smallest special subvariety of $G$ containing $\pi_{G}(W)$. We say that $W$ is exemplary if there are no larger distinguished components $W^{\prime}\supset W$ such that $\pi_{G}(W^{\prime})\subset B$.
  • We say a distinguished component $W$ in $G$ is dependent if $G^{\prime}\subset G$ is strict. We note that the unique non-dependent exemplary component is $V$ itself.
Pila and Tsimerman [PT22, Theorem 1.1] were able to deal with the case where $S=Y(1)^{n}$ is the $n$-copy of a modular (or Shimura) curve and $G=E^{n}$ is the self-$n$-product of an elliptic curve for $n\geq 1$. We prove the following theorem, which is a generalization of their main theorem to arbitrary (connected) Shimura varieties and commutative groups. The theorem describes all algebraic dependence among $V$-images of $n$ special points for any $n\geq 1$.
@@
Let $S$ be a connected Shimura variety, $G$ be a commutative connected algebraic group. Suppose $V\subset S\times G$ is an irreducible subvariety that maps finitely to $S$. Then, there are only finitely many exemplary components in $V$.
@@
Let $S,G,V$ be as above. Suppose moreover that the projections of $V$ to $S$ and $G$ are both dominant. Fix $n\in\mathbb{N}$. There exists a proper Zariski closed subset $S^{\prime}\subset S^{n}$ such that for any $n$ points $s_{1},\dots,s_{n}\in S$ satisfying $(s_{1},\dots,s_{n})\not\in S^{\prime}$, any $V$-images $g_{1},\dots,g_{n}$ of $s_{1},\dots,s_{n}$ are algebraically independent.
@@
For $n\geq 1$, let $\boldsymbol{s}=(s_{1},\cdots,s_{n})$ be a special point in $S^{n}$ and let $\boldsymbol{g}=(g_{1},\cdots,g_{n})$ be a point in $G^{n}$ such that $(\boldsymbol{s},\boldsymbol{g})\in V^{n}$. It follows from the definition that if $(\boldsymbol{s},\boldsymbol{g})\in W$ for some dependent exemplary component then $g_{1},\cdots,g_{n}$ are algebraically dependent in $G$. Conversely, if $g_{1},\cdots,g_{n}$ are algebraically dependent in $G$, then $(\boldsymbol{s},\boldsymbol{g})$ is a distinguished point and hence contained in some dependent exemplary component. ∎
@@
Suppose that $S\subset\mathcal{A}_{g}$ for $g\leq 3$. For $n\geq 1$ there exists a positive integer $D=D(n,S,G,V)$ such that if $s_{1},\dots,s_{n}$ are $D$-independent, then any $V$-image $x_{1},\dots,x_{n}$ of $s_{1},\dots,s_{n}$ are algebraically independent in $G$.
Taking $V$ to be the graph of the function $f\colon Y(1)^{n}\times Y(1)^{n}\to\mathbb{C}^{n}$ given by
@@
For $n\geq 1$, there exists a positive integer $D=D(n)$ such that if $x_{1},\dots,x_{n},$ $y_{1},\dots,y_{n}\in Y(1)$ are singular moduli that are $D$-independent, then there does not exist $(a_{1},\dots,a_{n})\in\mathbb{Z}^{n}\backslash\{\boldsymbol{0}\}$ such that
@@
We prove also that on low dimensional abelian varieties, the intersection of any finite-rank subgroup of $G$ with the set of $V$-images of special points on $S$ is contained in a finite union of proper special subvarieties. This generalizes Pila and Tsimerman’s result [PT22, Corollary 1.4] to higher dimensional abelian varieties, and Buium and Poonen’s [BP09, Theorem 1.1] in a uniform way, depending only on the rank of the subgroup.
@@
Suppose $G$ is defined over a field $K$ of characteristic 00 and suppose the Shimura datum for $S$ has $G_{S}$ simple. Let $\Gamma\subset G(K)$ be a subgroup of rank $r$ and let $\Gamma^{\prime}$ be its division group. Then, there exist $N(r,S,G,V)$ proper special subvarieties of $S$ such that the set of special points of $S$ with a $V$-image in $\Gamma^{\prime}$ is contained in a union of these proper special subvarieties.
A similar argument to [PT22, Proposition 3.2] proves that Theorem 1.2 is a consequence of the Zilber–Pink Conjecture. Also, Theorem 1.6 may be thought as a weak version of the André–Oort–Mordell–Lang conjecture in the case of subgroups of finite rank formulated by Baldi in [Bal18, Conjecture 5.2]. The original formulation of the André–Oort–Mordell–Lang conjecture is due to Pink, which can be found in [Pin05].
@@ 1.2. Organization of the paper
The notions of Shimura varieties, special and weakly special subvarieties are explained in Section 2. We also prove a mixed version of the Ax–Schanuel theorem for the uniformization of the product $S\times G$ in the form needed for our main theorem in this section. In Section 3, we collect and prove some arithmetic estimates. The main theorem is proved in Section 4 and its applications and corollaries are proved in Section 5.
@@ 2. Shimura varieties and the mixed Ax–Schanuel
Let $S$ be an algebraic variety and $q\colon X\to S$ be a universal covering of $S$. The Ax–Schanuel theorem gives information about the bi-algebraic varieties, algebraic varieties $V\subset X$ such that $q(V)$ is also algebraic, for the transcendental covering map $q$. The rough statement of the Ax–Schanuel theorem is that varieties of $X\times S$ should have a proper intersection with the graph of $q$ unless the projection of $V$ to $S$ is contained within a bi-algebraic variety. These bi-algebraic varieties are precisely the weakly special subvarieties. We first give precise definitions of bi-algebraic varieties and results in the case when $S$ is a commutative algebraic group or a Shimura variety, and then prove the result when $S$ is a product of the two.
@@ 2.1. Shimura Varieties
Let $S:=\operatorname{Sh}_{K}(G_{S},X)$ be the (connected) Shimura variety associated with a connected Shimura datum $(G_{S},X)$, where $G$ is a semisimple group of adjoint type and $K$ is a compact open subgroup of $G(\mathbb{A}_{f}).$ More information about Shimura varieties can be found in Milne ([Mil05]).
@@
For any $x\in X$, let $\mathbf{M}:=\operatorname{MT}(x)$ be the Mumford-Tate group of $x$, which is defined as the smallest $\mathbb{Q}$-subgroup $H$ of $G$ such that $x$ factors through $H_{\mathbb{R}}$. Let $X_{\mathbf{M}}$ be the $\mathbf{M}(\mathbb{R})$-conjugacy class of $x$. Then the image of $X_{\mathbf{M}}$ in $S=\Gamma_{g}\backslash X$ is a special subvariety of $S$. It is not hard to see that every special subvariety of $S$ arises this way. The action of $\mathbf{M}(\mathbb{R})$ on $X_{\mathbf{M}}$ factors through the adjoint group $\mathbf{M}^{\operatorname{ad}}(\mathbb{R})$, which is a direct product of its $\mathbb{Q}$-simple factors. Therefore, we can write $X_{\mathbf{M}}$ as a product
$X_{\mathbf{M}}=X_{1}\times X_{2},$
corresponding to the action of $M^{\operatorname{ad}}(\mathbb{R})$. A weakly special subvariety is the image of the fiber $\{x_{1}\}\times X_{2}$ or $X_{1}\times\{x_{2}\}$ for any $x_{1}\in X_{1}$ or $x_{2}\in X_{2}$. Therefore, a weakly special subvariety of $S$ is a special subvariety if and only if it contains a special point. The weakly special subvarieties of $S$ are precisely those subvarieties that are totally geodesic in $S$ ([Moo98, Section 4]), and they are also precisely the bi-algebraic subvarieties under the uniformization map $q\colon X\to S$. More detailed information on weakly special subvarieties of Shimura varieties can be found in [UY11].
@@
Then, the projection of $U$ to $S$ is contained in a proper weakly special subvariety of $S$.
2.2. Commutative Algebraic Groups
Let $G$ be a connected commutative algebraic group of dimension $g$ defined over $\mathbb{C}$ and let $\exp\colon\mathbb{C}^{g}\to G$ be the exponential map from $\operatorname{Lie}G\cong\mathbb{C}^{g}$ to $G$.
@@ 2.2. Commutative Algebraic Groups
For each connected subgroup $B\subset G$ and closed point $p\in G(\mathbb{C})$, we say that the algebraic subvariety $B+p\subset G$ is a weakly special subvariety of $G$. We say that it is a special subvariety of $G$ if $p\in G_{\mathrm{tors}}$ is a torsion point.
The special points of $G$ are special subvarieties of dimension 00, and hence they are precisely the torsion points of $G$. As before, a weakly special subvariety of $G$ is special precisely when it contains a special point.
The Ax–Schanuel theorem for commutative groups is a statement about covering map given by the $\operatorname{Lie}$-exponential $\exp$.
@@
Let $V\subset\mathbb{C}^{g}\times G$ be an algebraic subvariety and let $D\subset\mathbb{C}^{g}\times G$ be the graph of $\exp\colon\mathbb{C}^{g}\to G$. Let $U\subset V\cap D$ be an irreducible component such that
$\dim U>\dim V-\dim G.$
Then the projection of $U$ to $G$ is contained in a proper weakly special subvariety of $G$.
@@ 2.3. Mixed Ax–Schanuel
We will prove a version of the Ax–Schanuel theorem for the product uniformization map $q\times\exp\colon X\times\mathbb{C}^{g}\to S\times G$. This was proven by Pila and Tsimerman in [PT22] for the case when $S$ is a product of modular curves and $G$ is a product of elliptic curves. In that case, it was shown that the Ax–Schanuel theorem for the product follows from an Ax–Schanuel theorem for mixed Shimura varieties by Gao in [Gao20]. Using the same method, one can prove that the Ax–Schanuel theorem holds for the product $S\times G$ whenever $S$ is a Shimura variety of abelian type and $G$ is an abelian variety. However, to prove it for general Shimura varieties and general commutative groups, we will need stronger machinery from Blázquez-Sanz, Casale, Freitag, and Nagloo ([BSCFN23]). The translation into the language of (weakly) special subvarieties is given by Chiu ([Chi22]).
@@
Let $S$ be a Shimura variety and $G$ a connected commutative algebraic group. We say that an algebraic subvariety $V\subset S\times G$ is a weakly special subvariety of $S\times G$ if there exist weakly special subvarieties $S_{H}\subset S$ and $B\subset G$ such that $V=S_{H}\times B$. We say that $V$ is a special subvariety if $S_{H},B$ are both special.
@@
Let $V\subset(X\times\mathbb{C}^{g})\times(S\times G)$ be an algebraic subvariety and let $D\subset(X\times\mathbb{C}^{g})\times(S\times G)$ be the graph of $q\times\exp\colon(X\times\mathbb{C}^{g})\to(S\times G)$. Let $U\subset V\cap D$ be an irreducible component such that
$\dim U>\dim V-\dim(S\times G).$
Then the projection of $U$ to $S\times G$ is contained in a proper weakly special subvariety of $S\times G$.
@@
Let the Shimura datum of $S$ be $(G_{S},X)$. We may take $S$ to be a connected Shimura variety and $G$ to be a derived group. Applying [BSCFN23, Thm. 3.6] to $D\subset(X\times\mathbb{C}^{g})\times(S\times G)$ gives that the projection of $U$ in $S\times G$ is contained in a proper subvariety whose Galois group is a strict algebraic subgroup $H\subset G_{S}(\mathbb{C})\times\mathbb{G}_{a}(\mathbb{C})^{g}$. Since $G$ is a derived group, it has no abelian quotients and hence Goursat’s lemma says that the projection of $H$ to $G(\mathbb{C})$ or $\mathbb{G}_{a}(\mathbb{C})^{g}$ is not onto. By [Chi22, Thm. 3.2] for the Shimura variety side and [Ax72, Thm. 3] for the commutative group side, that means the projection of $U$ to $S$ or $G$ is contained in a proper weakly special subvariety. ∎